OpenStudy (anonymous):

a garden is shaped like a rectangle whose perimeter is 56 ft. the length is 2ft more than the width. find the width and length?

mathslover (mathslover):

let the width be x then length = 2 + x correct?

OpenStudy (anonymous):

Na is asking me for the length and width .

OpenStudy (anonymous):

i have to find wats the width and length.

OpenStudy (anonymous):

right.. so let x = width thus x + 2 would = length. i think he was confirming

mathslover (mathslover):

yes ... length = x+2 width = x perimeter = 56 since perimeter of a rectangle = 2(l+b) 2(x+2+x)=56 2(2x+2)=56 2x+2= 56/2 2x+2=28 2x=28-2 2x=26 x=13 length = x+2=13+2=15 width = x = 13

mathslover (mathslover):

got it?

OpenStudy (anonymous):

oddly enough i just divided 56 by four and subtracted one and added 2 lol. :)

OpenStudy (anonymous):

yea thats it i was doing the samething but it wouldnt work ......... i was a bit confused by it let me check again.

mathslover (mathslover):

:P ... that's what i did (indirectly) : \[\large{2(2x+2)=56}\] \[\large{2(2)(x+1)=56}\] \[\large{4(x+1)=56}\] \[\large{x+1=\frac{56}{4}}\] \[\large{x+1=14}\] \[\large{14-1=x}\] \[\large{x=13}\]

OpenStudy (anonymous):

yea so 13 and 15 are right thats wat it was asking for but just wants me to detail how i got it as well.

OpenStudy (anonymous):

id do it your way too :), with algebra

OpenStudy (anonymous):

i can get half way i just get stuck at times with the whole answers at times.

OpenStudy (anonymous):

thx guys.

OpenStudy (anonymous):

i divide to but i thought i was doing it the wrong way so i stop now that i see how u did it.